OpenStudy (anonymous):

what is the solution to this equation -6x+3=21

OpenStudy (shamim):

-6x+3=21 -6x=21-3 -6x=18 6x=-18 x=-3

OpenStudy (austinl):

-6x+3=21 Subtract 3 from both sides. (-6x+3)-3=21-3 -6x=18 Divide both sides by -6 (-6x)/6=18/6 x=-3